OpenStudy (fairy733189):

The area of a rectangular piece of land is 240 square meters. If the length of the land was 5 meters less and the width was 3 meters more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. Part B: What is the width of the land in meters? Show the steps of your work.

OpenStudy (fairy733189):

@satellite73

satellite73 (satellite73):

we call the width x, then say the length is y, we get the areas \(xy=240\)

satellite73 (satellite73):

If the length of the land was 5 meters less and the width was 3 meters more, the shape of the land would be a square. i guess that means what? \(x-5=y+3\) ? check it

satellite73 (satellite73):

yeah looks reasonalbe seem right to you?

OpenStudy (fairy733189):

It looks pretty right to me but what would be the answer and how do you solve?

satellite73 (satellite73):

we can solve \[x-5=y+3\] for say \(y\) by subtracting \(3\) from both sides, giveing \[x-8=y\]

satellite73 (satellite73):

then since the area is \(xy=240\) replace \(y\) by \(x-8\) and get the equation in one variable, namely \[x(x-8)=240\] solve that for \(x\)

OpenStudy (fairy733189):

The answer would be 124 then?

satellite73 (satellite73):

idk i didn't do it you want me to check?

satellite73 (satellite73):

seems unlikely

OpenStudy (fairy733189):

yes can you check please?

satellite73 (satellite73):

ok no

satellite73 (satellite73):

it is clearly not right \[124(124-8)\neq 240\] without doing the math

satellite73 (satellite73):

\[x(x-8)=240\] is a quadratic equation you gotta multiplicity out first \[x^2-8x=240\] is a start

OpenStudy (fairy733189):

so would I factor?

satellite73 (satellite73):

after you subtract \(240\) and write \[x^2-8x-240=0\] if you can factor, yes

satellite73 (satellite73):

i suck at factoring, if it was me, i would cheat

OpenStudy (fairy733189):

haha :D it's ok I know how to factor

satellite73 (satellite73):

ok let me know what you get

OpenStudy (fairy733189):

cant you also use the quadratic formula?

satellite73 (satellite73):

sure if you want this one has been cooked up to factor

satellite73 (satellite73):

since \(240=12\times 20\) you can use \[(x-20)(x+12)=0\]

OpenStudy (fairy733189):

ok and what would I write for my answer?

satellite73 (satellite73):

since if \((x+12)(x-20)=0\) you have \(x=-12\) or \(x=20\) and since the side cannot be negative, you know \(x=20\) is the solution

OpenStudy (fairy733189):

That makes complete sense, I understand! Thanks so much again!!!!

satellite73 (satellite73):

yw (again)

OpenStudy (fairy733189):

wait wouldn't the solution be 12 not 20 since the negative number is 20?

satellite73 (satellite73):

\[x-20=0\\ x=20\]

OpenStudy (fairy733189):

sorry im still a bit confused

satellite73 (satellite73):

we factored it as \[(x+12)(x-20)=0\] and need to solve for \(x\)

satellite73 (satellite73):

set each factor equal to zero and solve \[x+12=0\] subtract 12 get \[x=-12\] or \[x-20=0\] add 20 get \[x=20\]

OpenStudy (fairy733189):

oh Wow I understand it now :D Thanks AGAIN

satellite73 (satellite73):

yw

OpenStudy (fairy733189):

wait sorry again but what if it was (x - 12)(x + 20) how do you know which one is negative and which one to put as positive?

satellite73 (satellite73):

it wasn't

satellite73 (satellite73):

it was \[x^2-8x-240=(x+12)(x-20)\]

OpenStudy (fairy733189):

but how do you know which one will be negative and which one will be positive?

satellite73 (satellite73):

ooh because you start with \[x^2-8x-240\] right?

OpenStudy (fairy733189):

right

satellite73 (satellite73):

two numbers that multiply to give \(-240\) and add to \(-8\) so \(+12-20=-8\) and \(12\times (-20)=-240\) works

satellite73 (satellite73):

in other words, if you multiply out \[(x+12)(x-20)\] you get \[x^2-8x-240\]

OpenStudy (fairy733189):

yes I remember that now I don't know why but I completely forgot about that Thanks AGAIN im pretty sure this is the last question I needed to ask!

satellite73 (satellite73):

hey, keep 'em coming

OpenStudy (fairy733189):

haha :D

OpenStudy (fairy733189):

Well I guess i'm going to close this question now but thanks once again!